AS you hopefully have notice today was an access day, which the people present at RHIC spend fixing several problems. a) The CA-D engineers got the BFS into movement operation and it has been placed at 4 deg for the enxt measurements. b) Ramiro, Kris an myself looked into the TPM2 problem, and replaced (or rather moved) a couple of FEE cards around to optimize the response. c) Surveys were done for TFW2, so this should now be complete. d) Towards the very end a problem with FFS1440 caused us considerable grief; a card was exchanged just before close-up and then the whole MainFrame did not work. Eventuallt we got an access and for now it seems ok. e) a bad C4 channel was identified. f) the ZDC HV was lowered by ~*2; this has reduced most of the saturation and double pulsing problems, but will still have to be worked on. g) yesterday Dana recabled the towf,tfw2 readout to optimize the readout in terms of tdc modules etc. This should make future changes due to failing channels easier to accomodate. h) The repair of the spare L 1458 is proceeding. I may have left out a few items; it was a productive day overall. During tomorrow I will update the shift lists for the coming 2 weeks, taking into account the planned CA-D beam improvements. regards Flemming ---------------------------------------------------------------- Flemming Videbaek Physics Department Brookhaven National Laboratory e-mail: videbaek@bnl.gov phone: 631-344-4106 _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Wed Jan 7 19:24:06 2004
This archive was generated by hypermail 2.1.8 : Wed Jan 07 2004 - 19:24:24 EST